In an a.c. circuit I and E are given by I=100sin(100t+π⁄3) mA and E=100sin100t volt. what is the power dissipated in the circuit?
1
Expert's answer
2018-08-27T09:42:08-0400
P=(E_max·I_max·cos⁡〖π/3〗)/2=(100·100·0.5)/2=2500 W